

********************************************************************
UARS WEEKLY STATUS REPORT - 08 SEP 2000
********************************************************************
***********
SOLSTICE
***********
SOLSTICE is in good health, collecting solar data on all available orbital
tracking periods and stellar calibration observations whenever possible.
Our solar data is not affected by the larger than normal pointing errors
(2-4 arcminutes across a track) since they are within our calibrations.
Additionally, we continue to gather enough data to meet all the SOLSTICE
science objectives during this period of reduced tape recorder operations.
Level2 processing is halted due to bugs in the OA_SSPP_ATT service. CDHF
and CSC personnel are working on the problem.
***********
SUSIM
***********
SUSIM successfully conducted its monthly deuterium-lamp calibration scans
this week. Otherwise, the instrument continued to gather a normal set of
daily solar scans during the periods of scheduled telemetry.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
The PEM zenith spectrometers are activated in daylight mode.
At the present time, PEM energy deposition is not being generated.
Recent data configuration changes have affected PEM software causing multiple
failures. These failures are being reviewed by the PEM team. In order to
correct for new telemetry requirements, PEM will need to update processing
code with a patch delivery. The delivery patch may include changes which
would allow the PEM software to execute on a DEC Alpha machine. GSFC is
helping PEM with some DEC Alpha particulars which have caused PEM level 3
code to fail on DEC Alpha machines.
Updated HEPS calibration parameters have been finalized and are ready for
delivery. Updated MEPS quality calibration files have been finalized and are
ready for delivery. Level 3 software errors, discovered by Dr. Dean Pesnell,
have been fixed; his suggestions have been incorporated with in the PEM
production code. Both updated HEPS and MEPS calibration data, along with the
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
HALOE continues successful operations with all engineering parameters
within nominal operating region. Science data result continue to be
excellent. Use of the tape recorder for recording HALOE events generally
provides two thirds of the captured events while TDRS coverage provides the
remaining third.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
WINDII is next scheduled to operate on September 7 to 11.
***********
MPG
***********
The MPG participated in two meetings this week, both on
September 6. The first meeting was held to report the
status of FDF and FOT efforts regarding calibration of
Fixed Head Star Tracker 1. The second meeting was the
UARS Monthly Status Review meeting.
Processing of instrument Daily Activity Plans continues with no
problems. We are continuing to coordinate with the Flight
Operations Team and to assist the Project Scientist in the
scheduling of instrument operations.
***********
CDHF
***********
STATUS
- The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
- No reprocessing requested this reporting period.
- Notified the UKMO folks regarding day gaps and FILE_SIZE
differences on the current system. As of this reporting period, the
day gaps have been received and days have been re-ingested to correct
the FILE_SIZE difference.
- Data created in the month of July, 2000 will be previewed and loaded
into the Daac Transfer Tables September, 2000.
UARS CDHF REHOST:
Exercising the new MSS/SILO interface. Validating that it handles
stage/destage/recovery/exception events like current MSS/Filetek under
a production processing environment.
REHOST ISSUES:
Need Positive Response from all PI End Users about results from
August parallel test, prior to transition.
***********
FOT
***********
REPORTING PERIOD
This report covers 26 August 2000 (Orbit 48962, GMT Day 239) through 01
September 2000 (Orbit 49066, GMT Day 245).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 239/0053 48962 AXIS On
The beta angle ranged between the angles of 39.5 degrees to 48.3 degrees for
this report period. The beta angle is now decreasing toward a minimum angle of
+0 degrees on 18 September 2000 (DOY 262).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 0.8 to 1.1 degrees C. Additional battery performance data for each
battery on 25 August 2000 and 01 September 2000 are provided in Appendix D.
The clock error ranged from -11.0 to 11.0 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
¥ OUT-OF-LIMITS = 00
¥ ALERTS = 01
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
¥ FOT = 43
¥ PEM = 03
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
¥ FOT = 00
Three (3) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 00/153 - 2287.496961
· Transponder B: 00/213 - 2287.499288
GROUND SYSTEM OPERATION
No operational support problem occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/244 show the observatory data loss to be 298 hours, 12
minutes, 51.897 seconds (an increase of 1 minute 28.528 seconds since the last
report period). This is a 0.39 percent data loss which equals a 99.61 percent
data capture for the mission.
The increase of 1 minute 28.528 seconds occurred over one (1) day:
· 01 minutes 28.528 seconds on Day 239 due to data hits.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT continues to operate FHST-1 for approximately one orbit each day. By
cycling the tracker on and off for such a short duration the degradation seems
to have been reduced. Scale factor calibrations are being provided on a
regular basis from FDF and are being uplinked when received. It appears that
the FOT has been able to stabilize the performance of FHST-1 for now and
continues to work on establishing another viable attitude control solution
should the tracker fail.
MEETINGS
The FOT attended a meeting to discuss FHST-1 on 28 August 2000.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for September 17, 2000 (DOY 261). This
maneuver will be a Reverse to Forward Flight maneuver (Southern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 01 Sept 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
48962 239/0053 00-365 PEM Power ON, AXIS 1 & 2
48965 239/0604 00-436 NBTR NOOPS D-239
48971 239/1621 00-422 FHST1 Power ON
48972 239/1803 00-422 FHST1 Power OFF
48972 239/1803 00-438 NBTR NOOPS D-240
48981 240/0744 00-438 NBTR NOOPS D-240
48986 240/1623 00-422 FHST1 Power ON
48986 240/1623 00-438 NBTR NOOPS D-240
48988 240/1848 00-422 FHST1 Power OFF
48994 241/0504 00-439 NBTR NOOPS D-241
48997 241/1010 00-440 OBC Clk Adj 0.00 to 20.6 msec/day
49001 241/1625 00-422 FHST1 Power ON
49001 241/1625 00-439 NBTR NOOPS D-241
49002 241/1808 00-441 PEM CAL - START
49002 241/1808 00-442 NBTR NOOPS D-242
49002 241/1808 00-422 FHST1 Power OFF
49005 241/2130 00-441 PEM CAL - STOP
49005 241/2130 00-443 FHST1 SCALE FACTOR UPDATE
49010 242/0609 00-442 NBTR NOOPS D-242
49016 242/1525 00-444 FHST1 UPDATE MAGNETIC FIELD COMP COEF
49016 242/1525 00-422 FHST1 Power ON - MONITOR MODE
49017 242/1710 00-442 NBTR NOOPS D-242
49017 242/1810 00-445 NBTR NOOPS D-243
49019 242/2032 00-446 OBC GENASCII TABLE 15 LOAD
49019 242/2032 00-447 FHST1 ST1 TO UPDATE MODE
49020 242/2132 00-448 OBC BASETAB MONDEF
49020 242/2132 00-449 OBC GENASCII TABLE 15 LOAD
49020 242/2210 00-422 FHST1 Power OFF
49025 243/0649 00-445 NBTR NOOPS D-243
49025 243/0649 00-450 OBC Clk Adj 20.6 to 0.00 msec/day
49031 243/1629 00-422 FHST1 Power ON
49031 243/1629 00-445 NBTR NOOPS D-243
49032 243/1811 00-451 NBTR NOOPS D-244
49032 243/1811 00-422 FHST1 Power OFF
49033 243/1952 00-452 FHST1 SCALE FACTOR UPDATE
49046 244/1632 00-422 FHST1 Power ON
49046 244/1632 00-451 NBTR NOOPS D-244
49047 244/1813 00-453 NBTR NOOPS D-245
49047 244/1813 00-422 FHST1 Power OFF
49054 245/0512 00-453 NBTR NOOPS D-245
49059 245/1203 00-454 OBC Clk Adj 0.00 to 20.6 msec/day
49061 245/1533 00-422 FHST-1 Power ON
49061 245/1634 00-453 NBTR NOOPS D-245
49062 245/1716 00-455 FHST-1 Scale Factor Update
49062 245/1814 00-422 FHST-1 Power OFF
49062 245/1814 00-456 NBTR NOOPS D-246
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 01 Sept 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
-None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
OBSZ1OOL UFLTR Z1 Residual Exceeded 100 Arcsec 48987
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 01 Sept 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
48997 241/1012 (28 August) from 0.00 to 20.6 msec/day
49025 243/0653 (30 August) from 20.6 to 0.00 msec/day
49059 245/1213 (01 September) from 0.00 to 20.6 msec/day
___________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 01 Sept 2000
25 Aug 2000 - GMT Day 238 Beta = 39.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- --------- ----------- ------ ------- ------
1 +728.0/+728.0 -3.05/-3.99 +0.0/+0.0 14.4 0.000 0.0
2 +39.2/-67.2 +5.79/+4.51 +28.0/-15.2 24.5 1.030 21.4
3 +22.4/-50.4 +5.15/+3.87 +27.6/-15.6 24.5 1.030 21.3
01 Sept 2000 - GMT Day 245 Beta = 48.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ---------- --------- ---------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/+0.0 14.4 0.000 0.0
2 +44.8/-11.2 +5.15/+3.87 +26.0/-15.2 25.0 1.040 19.8
3 +22.4/-33.6 +4.19/+2.92 +25.6/-15.2 25.0 1.038 19.7
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---------- ----------
no changes V/T 5
_______________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 01 Sept 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ----
-032 FHST 48211 188/2242 N/A FHST 2 Anomaly
SEVNTFW00/35